# FLawN-T5: An Empirical Examination of Effective Instruction-Tuning Data Mixtures for Legal Reasoning

### v1

@@ 5 Related Work
Domain-specific pretraining, covering areas such as medicine, law, and science, significantly enhances Language Model performance on related tasks (Beltagy et al., 2019; Gu et al., 2021; Chalkidis et al., 2020). SciBERT (Beltagy et al., 2019), for instance, underwent pretraining on a combination of computer science and biomedical papers, exemplifying this approach in the scientific domain. Other models like PubMedBERT (Gu et al., 2021) and BioBERT (Lee et al., 2020), specifically pretrained on biomedical datasets, have shown improvements in medical NLP tasks (Huang et al., 2019).
@@ 5.1 Domain-specific Legal Pretraining
In the legal domain, models such as LegalBERT, pretrained on 12 GB of English legal texts, demonstrated notable success in domain-specific challenges (Chalkidis et al., 2020). CaseLaw-BERT capitalized on the English Harvard Law case corpus spanning from 1965 to 2021 (Zheng et al., 2022), while Niklaus & Giofré (2022) pretrained LongFormer models on the Pile-of-Law (Henderson et al., 2022) using the replaced token detection task (Clark et al., 2020) for enhanced performance. Further advancements were made by Chalkidis et al. (2023), who developed new English legal LMs yielding superior results on LexFiles, a compilation of 11 sub-corpora from six English-speaking legal systems encompassing 19B tokens. Additionally, Niklaus et al. (2023b) introduced a vast multilingual legal corpus, training both monolingual and multilingual legal models to achieve state-of-the-art results on LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023a). Models have also been developed for specific jurisdictions, including the Swiss (Rasiah et al., 2023), Italian (Licari & Comandè, 2022), Romanian (Masala et al., 2021), and Spanish (Gutiérrez-Fandiño et al., 2021) legal systems. Despite the prevalence of smaller encoder-based legal-specific LMs, larger generative models in this space remain scarce. This work seeks to bridge that gap.
@@ 6 Conclusion and Future Work
In the future, we would like to extend LawInstruct with more high-quality datasets released after our experiments such as Long-form Legal Question Answering (Louis et al., 2023), Keyphrase Generation (Salaün et al., 2024), Negation Scope Resolution (Christen et al., 2023), or Legal Violation Detection (Bernsohn et al., 2024). Additionally, future work could work with legal professionals to identify and annotate new tasks for the legal domain. Future work could explore the benefits of synthetic data generation in the legal domain, either by generating responses from instructions (Wang et al., 2023c) or reversely, by generating instructions to responses (Köksal et al., 2024), but should take care to do detailed quality checks since hallucinated content may hurt more than improve, especially in the legal domain. Finally, it would be interesting to investigate overlap between the T5 pretraining dataset C4 and the MultiLegalPile to get a better understanding of the potential benefits of continued pretraining.
@@ Acknowledgements
We would like to thank Baq Haidri for help with debugging and John Kirchenbauer, Sang Michael Xie and David Hall for feedback on the project. Thanks to Neel Guha for help with debugging LegalBench and for feedback on analyzing results.
Appendix A Detailed Dataset Description
@@
Dataset License Languages Jurisdiction Tasks Subtask Examples Benchmark for Understanding Indian Legal Documents (BUILD) (Kalamkar et al., 2022) Unknown en India Text classification Rhetorical role 28,986 Brazilian Bar Exam (Delfino et al., 2017) Unknown pt Brazil Question answering Bar exam questions 2,130 Brazilian Court Decisions (Lage-Freitas et al., 2022) Unknown pt Brazil Text classification Judgment 3,234 Brazilian Court Decisions (Lage-Freitas et al., 2022) Unknown pt Brazil Text classification Decision Unanimity 1,715 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Multiple choice Judgment 1,225,922 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Judgment 612,961 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Area of law 612,961 BrCAD5 (Jacob de Menezes-Neto & Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Topic 1,838,883 BVADecisions (Walker et al., 2019) MIT en USA Text classification Rhetorical role 8,818 BVADecisions (Walker et al., 2019) MIT en USA Question answering Relevant rules 2 CAIL 2019 (Xiao et al., 2019) Unknown zh China Question answering Chinese legal case questions 39,333 CAIL 2022 (CAIL, 2022) Unknown zh China Text classification Charge/crime 10,448 CAIL 2022 (CAIL, 2022) Unknown zh China Argument & counter-argument 5,224 CAIL 2022 (CAIL, 2022) Unknown zh China Question answering Response to argument 5,224 Case Briefs (Case briefs, ) CC BY-NC en USA Question answering Legal analysis of facts 2,619 CaseHOLD (Zheng et al., 2021) CC-BY en USA Multiple choice Legal holding statements 45,000 Change My View (Tan et al., 2016) Unknown en N/A Argument & counter-argument 3,456 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Question generation Entailed question 1,774 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Natural language inference Passage entailment 125,954 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Question answering Relevant legal rule 1,774 ContractNLI (Koreeda & Manning, 2021) CC BY-NC en USA Natural language inference Premise hypothesis entailment 14,010 COVID-19 Emergency Measures (EXCEPTIUS) (Tziafas et al., 2021) Unknown en, fr, hu, it, nb, nl, pl EU Text classification Measure type 3,312 European Court of Human Rights (ECtHR) (Chalkidis et al., 2021b) CC BY-NC-SA 4.0 en EU Text classification (multi-label) Violated article 9,000 European Court of Human Rights (ECtHR) (Chalkidis et al., 2021b) CC BY-NC-SA 4.0 en EU Text classification (multi-label) Allegedly violated article 9,000 EOIR (Henderson et al., 2022) CC BY-NC-SA 4.0 en USA Text classification Pseudonymity 8,089 EURLEX (Chalkidis et al., 2019) CC BY-SA 4.0 en EU Text classification EuroVoc core concepts 55,000 EUR-Lex-Sum (Aumiller et al., 2022) CC BY 4.0 24 EU langs EU Summarization EU Legal Acts 22,989 German Argument Mining (Urchs et al., 2021) CC BY 4.0 de Germany Text classification Argumentative function 19,271 German Rental Agreements (Steinberger et al., 2006) Unknown de Germany Text classification Semantic type 3,292 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Volume (coarse thematic topic) 28,536 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Chapter (intermediate thematic topic) 28,536 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Subject (fine-grain thematic topic) 28,536 Greek Legal NER (elNER) (Bartziokas et al., 2020) CC BY-NC-SA 4.0 el Greece Named entity recognition Greek legal entities 17,699 ILDC (Malik et al., 2021) CC BY-NC en India Text classification Judgment 37,387 International Citizenship Law (Vink et al., 2021) CC BY 4.0 en International Question answering Citizenship acquisition 6,460 International Citizenship Law (Vink et al., 2021) CC BY 4.0 en International Question answering Citizenship loss 2,850 JEC-QA (Zhong et al., 2020) CC BY-NC-ND zh China Multiple choice National Judicial Examination of China 21,072 Korean Legal QA (Jeon, 2021) Academic use only ko South Korea Question answering Relevant law 1,830 LawngNLI (Bruno & Roth, 2022) MIT en USA Natural language inference Premise hypothesis entailment 1,142,304 LBOX OPEN (Hwang et al., 2022) CC BY-NC ko South Korea Text classification Judgment 12,142 LBOX OPEN (Hwang et al., 2022) CC BY-NC ko South Korea Text classification Relevant statutes 13,317 LEDGAR (Tuggener et al., 2020) CC BY-NC en USA Text classification Contract provision category 60,000 Legal Case Document Summarization (Shukla et al., 2022; Bhattacharya et al., 2019) CC BY-SA en India Summarization Indian Supreme Court 7,080 Legal Case Summarization (Shukla et al., 2022; Bhattacharya et al., 2019) CC BY-SA en UK Summarization UK Supreme Court 693 LegalNERo (Pais et al., 2021) CC0 1.0 ro Romania Named entity recognition Romanian legal entities 7,552 LegalQA (LegalQA, ) Unknown zh China Question answering Legal advice 21,946 LeNER-Br (Luz de Araujo et al., 2018) Unknown pt Brazil Named entity recognition Brazilian legal entities 7,828 Littleton (Basu et al., 2022) MIT en USA Question answering Relevant future interests 131 Littleton (Basu et al., 2022) MIT en USA Question answering Event graph 143 MAPA (de Gibert Bonet et al., 2022) CC BY-NC 4.0 24 EU langs EU Named entity recognition Coarse-grained 27,823 MAPA (de Gibert Bonet et al., 2022) CC BY-NC 4.0 24 EU langs EU Named entity recognition Fine-grained 27,823 MAUD (Wang et al., 2023a) CC BY en USA Multiple choice Merger agreement questions 10,751 MAUD (Wang et al., 2023a) CC BY en USA Text classification Deal point category 25,827 MAUD (Wang et al., 2023a) CC BY en USA Text classification Question type 25,827 MAUD (Wang et al., 2023a) CC BY en USA Text classification Text type 25,827 Mining Legal Arguments (Habernal et al., 2022) Apache-2.0 en EU Named entity recognition Actors 31,852 Mining Legal Arguments (Habernal et al., 2022) Apache-2.0 en EU Named entity recognition Argument type 31,852 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (coarse level) 1,265,000 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (intermediate level) 911,798 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (fine-grain level) 1,265,000 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Long to short 2,210 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Long to tiny 1,130 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Short to tiny 1,129 Natural Instructions (BillSum) (Kornilova & Eidelman, 2019) CC0 1.0 en USA Summarization U.S Congressional and California state bills 25,200 Natural Instructions (CAIL 2018) (Xiao et al., 2018) Unknown zh China Question answering Judgment 5,988 Natural Instructions (CaseHOLD) (Zheng et al., 2021) CC-BY en USA Multiple choice Correct answer 5,988 Natural Instructions (CaseHOLD) (Zheng et al., 2021) CC-BY en USA Multiple choice Incorrect answer 5,988 Natural Instructions (CUAD) (Hendrycks et al., 2021c) CC BY 4.0 en Question answering Information relevant for contract review 2,442 Natural Instructions (CUAD) (Hendrycks et al., 2021c) CC BY 4.0 en USA Question generation Questions relevant for contract review 2,442 Natural Instructions (EURLEX) (Chalkidis et al., 2019) CC BY-SA 4.0 en EU Text classification Regulation, decisions, or directive 5,850 Natural Instructions (EURLEX) (Aumiller et al., 2022) CC BY-SA 4.0 en EU Summarization EU Legal Acts 3,900 Natural Instructions (OPP-115) (Wilson et al., 2016) CC BY-NC en USA Question answering Type of information used by website 18,480 Natural Instructions (OPP-115) (Wilson et al., 2016) CC BY-NC en USA Question answering Purpose of privacy policy 18,474 Natural Instructions (Overruling) (Zheng et al., 2021) Unknown en USA Text classification Sentence is overruling 14,370 OLC Memos (Henderson et al., 2022) CC BY-NC en USA Question answering Write a legal research memo 1,038 Online ToS (Drawzeski et al., 2021) CC BY-NC 2.5 de, en, it, pt Unknown Text classification Clause topic 19,942 Online ToS (Drawzeski et al., 2021) CC BY-NC 2.5 de, en, it, pt Unknown Text classification Unfair contractual term type 2,074 Plain English Contracts Summarization (Manor & Li, 2019) Unknown en USA Summarization Software licenses, ToS 446 PrivacyQA (Ravichander et al., 2019) MIT en Unknown Question answering Contents of privacy policies 185,200 PrivacySummarization (Keymanesh et al., 2020) MIT en USA Summarization Privacy policies, ToS, and cookie policies 5,751 RedditLegalQA (Henderson et al., 2022) CC BY 4.0 en Unknown Question answering Legal advice from r/legaladvice 192,953 Sara (Holzenberger et al., 2020) Unknown en USA Natural language entailment Fact entailment 176 Sara (Holzenberger et al., 2020) Unknown en USA Question answering Tax liability 160 SaraProlog (Holzenberger et al., 2020) Unknown en USA Question answering Fact pattern to prolog code 376 SaraProlog (Holzenberger et al., 2020) Unknown en USA Question answering Tax statute to prolog code 9 Short Answer Feedback (Filighera et al., 2022) CC BY 4.0 de Germany Question answering Answer question about German law 1,596 Short Answer Feedback (Filighera et al., 2022) CC BY 4.0 de Germany Question answering Feedback rating for answer 1,596 Spanish Labor Law (Calleja et al., 2021) CC BY 4.0 es Spain Extractive question answering Answer question about Spanish labor law 111 StackExchange Questions (Law) (Law Stack Exchange, ) CC BY-SA en Unknown Question answering Online legal forum 10,158 The Supreme Court Database (Spaeth et al., 2020) CC BY-NC 3.0 en USA Text classification Issue areas 5,000 Swiss Federal Supreme Court (Rasiah et al., 2023) CC BY 4.0 de, fr Text generation Case considerations sections (lower court) 26 Swiss Courts (Rasiah et al., 2023) CC BY 4.0 de, fr, it Switzerland Text generation Case considerations sections (same court) 234,313 Swiss Federal Supreme Court (Rasiah et al., 2023) CC BY 4.0 de, fr, it Switzerland Text classification Case criticality (based on citations) 91,075 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Multiple choice Judgment 477,636 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it Switzerland Text classification Judgment 385,719 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Text classification Area of law 18,162 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Text classification Subarea of law 18,162 Swiss Federal Supreme Court (Leading Decisions) (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it Switzerland Text classification Location (canton, region) 42,342 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, fr, it, rm Switzerland Text classification Abbreviation 11,045 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Canton 35,698 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Short description 3,747 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Title 35,359 Thai Supreme Court Cases (TSCC)(Thanh et al., 2021) Academic use only th Thailand Question answering Relevant legal articles (Thai Criminal Code) 2,883 Turkish Constitutional Court (Mumcuoğlu et al., 2021) CC BY 4.0 tr Turkey Multiple choice Judgment 1,804 Turkish Constitutional Court (Mumcuoğlu et al., 2021) CC BY 4.0 tr Turkey Text classification Judgment 902 Unfair ToS (Lippi et al., 2019) Unknown en USA Text classification (multi-label) Unfair contractual term type 5,532 U.S Class Actions (Semo et al., 2022) GPL-3.0 en USA Text classification Judgment 3,000 Valid Wills (Kwak et al., 2022) Unknown en USA Text classification Statement supported by law/condition 1,512
@@
Appendix B Detailed Experimental Setup
B.1 Inexplicable Behaviour at the XXL size

### v2

@@ 5 Related Work
Domain-specific pretraining, covering areas such as medicine, law, and science, significantly enhances Language Model performance on related tasks (Beltagy et al., 2019; Gu et al., 2021; Chalkidis et al., 2020). SciBERT (Beltagy et al., 2019), for instance, was pretrained on a mix of computer science and biomedical papers, exemplifying this approach in the scientific domain. Other models like PubMedBERT (Gu et al., 2021) and BioBERT (Lee et al., 2020), specifically pretrained on biomedical datasets, have shown improvements in medical NLP tasks (Huang et al., 2019).
@@ 5.1 Domain-specific Legal Pretraining
In the legal domain, models such as LegalBERT, pretrained on 12 GB of English legal texts, demonstrated notable success in domain-specific challenges (Chalkidis et al., 2020). CaseLaw-BERT capitalized on the English Harvard Law case corpus spanning from 1965 to 2021 (Zheng et al., 2022), while Niklaus and Giofré (2022) pretrained LongFormer models on the Pile-of-Law (Henderson et al., 2022) using the replaced token detection task (Clark et al., 2020) for enhanced performance. Further advancements were made by Chalkidis et al. (2023), who developed new English legal LMs yielding superior results on LexFiles, a compilation of 11 sub-corpora from six English-speaking legal systems encompassing 19B tokens. Additionally, Niklaus et al. (2024) introduced a vast multilingual legal corpus, training both monolingual and multilingual legal models to achieve state-of-the-art results on LexGLUE (Chalkidis et al., 2022) and LEXTREME (Niklaus et al., 2023). Models have also been developed for specific jurisdictions, including the Swiss (Rasiah et al., 2023), Italian (Licari and Comandè, 2022), Romanian (Masala et al., 2021), and Spanish (Gutiérrez-Fandiño et al., 2021) legal systems. Despite the prevalence of smaller encoder-based legal-specific LMs, larger generative models in this space remain scarce. This work seeks to bridge that gap.
@@ 6 Conclusion
Limitations & Future Work
Our use of template-based instruction creation may restrict the variety of instructions, potentially affecting the model’s ability to handle more diverse or novel legal queries effectively. While we already tried to address this by paraphrasing the instructions with GPT-4, the diversity may still be limited. To alleviate this problem, we could create synthetic data either by generating responses from instructions (Wang et al., 2023c) or reversely, by generating instructions to responses (Köksal et al., 2024). It is important to take care to do detailed quality checks since hallucinated content may hurt more than improve, especially in the legal domain. Another way to alleviate this diversity problem is working with legal professionals to identify and annotate new tasks for the legal domain. However, this route is out of reach for many academic efforts due to large salaries of qualified lawyers.
It would be great to conduct a qualitative human expert study on the model outputs. Unfortunately, a rigorous human evaluation with qualified legal experts across multiple jurisdictions was beyond our current resources. As LegalBench and MMLU are multiple-choice QA benchmarks, the utility of human evaluation is limited compared to long-form generative tasks. The objective nature of these benchmarks allows for reliable automated evaluation. However, as future work expands to evaluating on more open ended generation tasks we definitely recommend conducting human evaluations.
To our surprise, continued pretraining only benefited at the Small model size, but not at larger sizes. Due to our focus on instruction tuning and limited budget, we were not able to study this effect in more detail. In future work, we would like to study the robustness of our findings across model sizes. We hypothesize that methods like mixing in data from the original training set, using smaller learning rates, and adding loss terms to discourage the weights to depart too much from the original model could potentially lead to different conclusions.
In the future, we would like to extend LawInstruct with more high-quality datasets released after our experiments such as Long-form Legal Question Answering (Louis et al., 2023), Keyphrase Generation (Salaün et al., 2024), Negation Scope Resolution (Christen et al., 2023), or Legal Violation Detection (Bernsohn et al., 2024).
Additionally, it would be interesting to investigate the overlap between the T5 pretraining dataset C4 and the MultiLegalPile to get a better understanding of the potential benefits of continued pretraining.
@@ Acknowledgements
We would like to thank Baq Haidri and Irhum Shafkat for help with debugging and John Kirchenbauer, Sang Michael Xie and David Hall for feedback on the project. Thanks to Neel Guha for help with debugging LegalBench and for feedback on analyzing results. Thanks to the Stanford NLP group and the anonymous reviewers for feedback on the draft.
Appendix A Use of AI Assistants
We used ChatGPT 3.5 and 4 for shortening texts and editing LaTeX more efficiently.
Appendix B Detailed Dataset Description
@@
Dataset License Languages Jurisdiction Tasks Subtask Examples Benchmark for Understanding Indian Legal Documents (BUILD) (Kalamkar et al., 2022) Unknown en India Text classification Rhetorical role 28,986 Brazilian Bar Exam (Delfino et al., 2017) Unknown pt Brazil Question answering Bar exam questions 2,130 Brazilian Court Decisions (Lage-Freitas et al., 2022) Unknown pt Brazil Text classification Judgment 3,234 Brazilian Court Decisions (Lage-Freitas et al., 2022) Unknown pt Brazil Text classification Decision Unanimity 1,715 BrCAD5 (Jacob de Menezes-Neto and Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Multiple choice Judgment 1,225,922 BrCAD5 (Jacob de Menezes-Neto and Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Judgment 612,961 BrCAD5 (Jacob de Menezes-Neto and Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Area of law 612,961 BrCAD5 (Jacob de Menezes-Neto and Clementino, 2022) CC BY-NC-SA 4.0 pt Brazil Text classification Topic 1,838,883 BVADecisions (Walker et al., 2019) MIT en USA Text classification Rhetorical role 8,818 BVADecisions (Walker et al., 2019) MIT en USA Question answering Relevant rules 2 CAIL 2019 (Xiao et al., 2019) Unknown zh China Question answering Chinese legal case questions 39,333 CAIL 2022 (CAIL, 2022) Unknown zh China Text classification Charge/crime 10,448 CAIL 2022 (CAIL, 2022) Unknown zh China Argument & counter-argument 5,224 CAIL 2022 (CAIL, 2022) Unknown zh China Question answering Response to argument 5,224 Case Briefs (Case briefs, ) CC BY-NC en USA Question answering Legal analysis of facts 2,619 CaseHOLD (Zheng et al., 2021) CC-BY en USA Multiple choice Legal holding statements 45,000 Change My View (Tan et al., 2016) Unknown en N/A Argument & counter-argument 3,456 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Question generation Entailed question 1,774 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Natural language inference Passage entailment 125,954 COLIEE (Kim et al., 2022) Academic use only en, jp Canada/Japan Question answering Relevant legal rule 1,774 ContractNLI (Koreeda and Manning, 2021) CC BY-NC en USA Natural language inference Premise hypothesis entailment 14,010 COVID-19 Emergency Measures (EXCEPTIUS) (Tziafas et al., 2021) Unknown en, fr, hu, it, nb, nl, pl EU Text classification Measure type 3,312 European Court of Human Rights (ECtHR) (Chalkidis et al., 2021b) CC BY-NC-SA 4.0 en EU Text classification (multi-label) Violated article 9,000 European Court of Human Rights (ECtHR) (Chalkidis et al., 2021b) CC BY-NC-SA 4.0 en EU Text classification (multi-label) Allegedly violated article 9,000 EOIR (Henderson et al., 2022) CC BY-NC-SA 4.0 en USA Text classification Pseudonymity 8,089 EURLEX (Chalkidis et al., 2019) CC BY-SA 4.0 en EU Text classification EuroVoc core concepts 55,000 EUR-Lex-Sum (Aumiller et al., 2022) CC BY 4.0 24 EU langs EU Summarization EU Legal Acts 22,989 German Argument Mining (Urchs et al., 2021) CC BY 4.0 de Germany Text classification Argumentative function 19,271 German Rental Agreements (Steinberger et al., 2006) Unknown de Germany Text classification Semantic type 3,292 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Volume (coarse thematic topic) 28,536 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Chapter (intermediate thematic topic) 28,536 Greek Legal Code (Papaloukas et al., 2021) CC BY 4.0 el Greece Text classification Subject (fine-grain thematic topic) 28,536 Greek Legal NER (elNER) (Bartziokas et al., 2020) CC BY-NC-SA 4.0 el Greece Named entity recognition Greek legal entities 17,699 ILDC (Malik et al., 2021) CC BY-NC en India Text classification Judgment 37,387 International Citizenship Law (Vink et al., 2021) CC BY 4.0 en International Question answering Citizenship acquisition 6,460 International Citizenship Law (Vink et al., 2021) CC BY 4.0 en International Question answering Citizenship loss 2,850 JEC-QA (Zhong et al., 2020) CC BY-NC-ND zh China Multiple choice National Judicial Examination of China 21,072 Korean Legal QA (Jeon, 2021) Academic use only ko South Korea Question answering Relevant law 1,830 LawngNLI (Bruno and Roth, 2022) MIT en USA Natural language inference Premise hypothesis entailment 1,142,304 LBOX OPEN (Hwang et al., 2022) CC BY-NC ko South Korea Text classification Judgment 12,142 LBOX OPEN (Hwang et al., 2022) CC BY-NC ko South Korea Text classification Relevant statutes 13,317 LEDGAR (Tuggener et al., 2020) CC BY-NC en USA Text classification Contract provision category 60,000 Legal Case Document Summarization (Shukla et al., 2022; Bhattacharya et al., 2019) CC BY-SA en India Summarization Indian Supreme Court 7,080 Legal Case Summarization (Shukla et al., 2022; Bhattacharya et al., 2019) CC BY-SA en UK Summarization UK Supreme Court 693 LegalNERo (Pais et al., 2021) CC0 1.0 ro Romania Named entity recognition Romanian legal entities 7,552 LegalQA (LegalQA, ) Unknown zh China Question answering Legal advice 21,946 LeNER-Br (Luz de Araujo et al., 2018) Unknown pt Brazil Named entity recognition Brazilian legal entities 7,828 Littleton (Basu et al., 2022) MIT en USA Question answering Relevant future interests 131 Littleton (Basu et al., 2022) MIT en USA Question answering Event graph 143 MAPA (de Gibert Bonet et al., 2022) CC BY-NC 4.0 24 EU langs EU Named entity recognition Coarse-grained 27,823 MAPA (de Gibert Bonet et al., 2022) CC BY-NC 4.0 24 EU langs EU Named entity recognition Fine-grained 27,823 MAUD (Wang et al., 2023a) CC BY en USA Multiple choice Merger agreement questions 10,751 MAUD (Wang et al., 2023a) CC BY en USA Text classification Deal point category 25,827 MAUD (Wang et al., 2023a) CC BY en USA Text classification Question type 25,827 MAUD (Wang et al., 2023a) CC BY en USA Text classification Text type 25,827 Mining Legal Arguments (Habernal et al., 2022) Apache-2.0 en EU Named entity recognition Actors 31,852 Mining Legal Arguments (Habernal et al., 2022) Apache-2.0 en EU Named entity recognition Argument type 31,852 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (coarse level) 1,265,000 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (intermediate level) 911,798 MultiEURLEX (Chalkidis et al., 2021a) CC BY-SA 24 EU langs EU Text classification (multi-label) EuroVoc taxonomy (fine-grain level) 1,265,000 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Long to short 2,210 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Long to tiny 1,130 Multi-LexSum (Shen et al., 2022) ODC-By en USA Summarization Short to tiny 1,129 Natural Instructions (BillSum) (Kornilova and Eidelman, 2019) CC0 1.0 en USA Summarization U.S Congressional and California state bills 25,200 Natural Instructions (CAIL 2018) (Xiao et al., 2018) Unknown zh China Question answering Judgment 5,988 Natural Instructions (CaseHOLD) (Zheng et al., 2021) CC-BY en USA Multiple choice Correct answer 5,988 Natural Instructions (CaseHOLD) (Zheng et al., 2021) CC-BY en USA Multiple choice Incorrect answer 5,988 Natural Instructions (CUAD) (Hendrycks et al., 2021c) CC BY 4.0 en Question answering Information relevant for contract review 2,442 Natural Instructions (CUAD) (Hendrycks et al., 2021c) CC BY 4.0 en USA Question generation Questions relevant for contract review 2,442 Natural Instructions (EURLEX) (Chalkidis et al., 2019) CC BY-SA 4.0 en EU Text classification Regulation, decisions, or directive 5,850 Natural Instructions (EURLEX) (Aumiller et al., 2022) CC BY-SA 4.0 en EU Summarization EU Legal Acts 3,900 Natural Instructions (OPP-115) (Wilson et al., 2016) CC BY-NC en USA Question answering Type of information used by website 18,480 Natural Instructions (OPP-115) (Wilson et al., 2016) CC BY-NC en USA Question answering Purpose of privacy policy 18,474 Natural Instructions (Overruling) (Zheng et al., 2021) Unknown en USA Text classification Sentence is overruling 14,370 OLC Memos (Henderson et al., 2022) CC BY-NC en USA Question answering Write a legal research memo 1,038 Online ToS (Drawzeski et al., 2021) CC BY-NC 2.5 de, en, it, pt Unknown Text classification Clause topic 19,942 Online ToS (Drawzeski et al., 2021) CC BY-NC 2.5 de, en, it, pt Unknown Text classification Unfair contractual term type 2,074 Plain English Contracts Summarization (Manor and Li, 2019) Unknown en USA Summarization Software licenses, ToS 446 PrivacyQA (Ravichander et al., 2019) MIT en Unknown Question answering Contents of privacy policies 185,200 PrivacySummarization (Keymanesh et al., 2020) MIT en USA Summarization Privacy policies, ToS, and cookie policies 5,751 RedditLegalQA (Henderson et al., 2022) CC BY 4.0 en Unknown Question answering Legal advice from r/legaladvice 192,953 Sara (Holzenberger et al., 2020) Unknown en USA Natural language entailment Fact entailment 176 Sara (Holzenberger et al., 2020) Unknown en USA Question answering Tax liability 160 SaraProlog (Holzenberger et al., 2020) Unknown en USA Question answering Fact pattern to prolog code 376 SaraProlog (Holzenberger et al., 2020) Unknown en USA Question answering Tax statute to prolog code 9 Short Answer Feedback (Filighera et al., 2022) CC BY 4.0 de Germany Question answering Answer question about German law 1,596 Short Answer Feedback (Filighera et al., 2022) CC BY 4.0 de Germany Question answering Feedback rating for answer 1,596 Spanish Labor Law (Calleja et al., 2021) CC BY 4.0 es Spain Extractive question answering Answer question about Spanish labor law 111 StackExchange Questions (Law) (Law Stack Exchange, ) CC BY-SA en Unknown Question answering Online legal forum 10,158 The Supreme Court Database (Spaeth et al., 2020) CC BY-NC 3.0 en USA Text classification Issue areas 5,000 Swiss Federal Supreme Court (Rasiah et al., 2023) CC BY 4.0 de, fr Text generation Case considerations sections (lower court) 26 Swiss Courts (Rasiah et al., 2023) CC BY 4.0 de, fr, it Switzerland Text generation Case considerations sections (same court) 234,313 Swiss Federal Supreme Court (Rasiah et al., 2023; Stern et al., 2024) CC BY 4.0 de, fr, it Switzerland Text classification Case criticality (based on citations) 91,075 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Multiple choice Judgment 477,636 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it Switzerland Text classification Judgment 385,719 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Text classification Area of law 18,162 Swiss Courts (Rasiah et al., 2023; Niklaus et al., 2021) CC BY 4.0 de, fr, it, en Switzerland Text classification Subarea of law 18,162 Swiss Federal Supreme Court (Leading Decisions) (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it Switzerland Text classification Location (canton, region) 42,342 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, fr, it, rm Switzerland Text classification Abbreviation 11,045 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Canton 35,698 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Short description 3,747 Swiss Legislation (Rasiah et al., 2023) CC BY 4.0 de, en, fr, it, rm Switzerland Text classification Title 35,359 Thai Supreme Court Cases (TSCC)(Thanh et al., 2021) Academic use only th Thailand Question answering Relevant legal articles (Thai Criminal Code) 2,883 Turkish Constitutional Court (Mumcuoğlu et al., 2021) CC BY 4.0 tr Turkey Multiple choice Judgment 1,804 Turkish Constitutional Court (Mumcuoğlu et al., 2021) CC BY 4.0 tr Turkey Text classification Judgment 902 Unfair ToS (Lippi et al., 2019) Unknown en USA Text classification (multi-label) Unfair contractual term type 5,532 U.S Class Actions (Semo et al., 2022) GPL-3.0 en USA Text classification Judgment 3,000 Valid Wills (Kwak et al., 2022) Unknown en USA Text classification Statement supported by law/condition 1,512
Appendix C Detailed Experimental Setup
C.1 Inexplicable Behaviour at the XXL Size